OpenStudy (anonymous):

differentiate \[1 \over 1+5x\]

OpenStudy (anonymous):

-5/(5x+1)^2

OpenStudy (anonymous):

How?

OpenStudy (hoblos):

1/(1+5x) = (1+5x)^(-1)

OpenStudy (anonymous):

Use the Quotient rule along with the product rule.

OpenStudy (anonymous):

Hmm, can you show the steps?

OpenStudy (hoblos):

you can use standard differentiation [(1+5x)^(-1)]' = -1[(1+5x)^(-2) * 5) = -5(5x+1)^(-2)

OpenStudy (anonymous):

Thanks :)

OpenStudy (hoblos):

my pleasure XD

sam (.sam.):